OpenStudy (anonymous):

Dora can run 3/5 mi in 3/20 h. How far can she run in 1 h?

OpenStudy (igreen):

3/5 = 0.6 3 / 20 = 0.15 0.6 / 0.15 = ?

OpenStudy (anonymous):

thx green

OpenStudy (igreen):

No problem.